Find the derivative of f(v)=3 sqrt v-2ve/v


1
Expert's answer
2021-05-24T16:19:47-0400

"f(x) =3\sqrt v- 2v"

"f(x) = 3(1\/2)(v^{-1\/2})- 2"

"f(x) = \dfrac{3}{2\sqrt v} - 2"